OpenStudy (anonymous):

solve 8cos^2(x)-6cos(x)+1=0

Nnesha (nnesha):

that's the quadratic equation let cos(x) = y to make it simple \[\huge\rm 8y^2-6y+1=0\] now can you factor the quadratic equation ?

Nnesha (nnesha):

A)yes B) no C) maybe pick one

OpenStudy (anonymous):

yes

Nnesha (nnesha):

okay :=)) good try to factor and then post your work

OpenStudy (anonymous):

8y^2 -6y +1 = (2y-1)(4y-1) ---> 2y*4y=8y^2, 2y*-1=-2y, -1*4y= -4y, -1*-1=1 therefore 8y^2 -2y -4y +1= 8y^2 -6y +1 so it is factorable and above is it's factored form

Nnesha (nnesha):

that's correct! nice job \[\rm \color{ReD}{(2y-1)}\color{blue}{(4y-1)}=0\] set them equal to zero \[\rm \color{reD}{2y-1}=0\] and \[\rm \color{blue}{4y-1}=0\] solve for y

Nnesha (nnesha):

btw we should multiply the same two numbers to get the product of AC (a=leading coefficient c=constant term)so instead multiplying separately 2y and 4y by -1 you can just do -2y * -4y = 8y^2 \[\rm negative*negative =positive \]

OpenStudy (anonymous):

alright I will keep that in mind thank you

Nnesha (nnesha):

bec \[\rm -2y*-4y=8y^2\] \[\rm -2y-6y=-6y\] \[2y+4y\cancel{= } -6y\]

Nnesha (nnesha):

but that's the same thing after we multiply them by negative one alright so what do you get for y ? :=))

OpenStudy (anonymous):

y= 1/2 , 1/4

Nnesha (nnesha):

that's correct now we can replace y with cos(x) so cos(x) =1/2 , cos(x)=1/4 to find an angle we should take invs cos here is an example \[\large\rm \sin(x) = a \rightarrow x=\sin^{-1}(a)\] invssin is same as arcsin invscos is same as arccos

Nnesha (nnesha):

^^in that example x is an angle

OpenStudy (anonymous):

so arccos(1/2) = X1, and arccos(1/4) = X2.

Nnesha (nnesha):

yes that's correct you can use unit circle for 1/2 but i guess you have to use calculator for 1/4 i hope you know on the unit circle x-coordinate = cos and y-coordinate = sin

OpenStudy (anonymous):

alirght so x = 1/3pi and x = 1.318...

Nnesha (nnesha):

oh do you need in radian or degree ? if you need radian then keep it in fraction form

Nnesha (nnesha):

 there is another angle where cos =1/2

OpenStudy (anonymous):

don't worry it is all good, huge help btw thank you!

Nnesha (nnesha):

np :=))